# Structural Properties of Entropic Vectors and Stability of the Ingleton Inequality

### v1

@@ 3.2 Proof of Lemma D
$=I(\tilde{A}_{k}:X,Y|X)$
/* since $\tilde{A}_{k}\to(XY)\to(A,B,C_{0},\ldots,C_{n-1})$ is a Markov chain */
@@
$H(W_{k}|X)=k\cdot\left(H(A|X)-H(A|X,Y)\right)=k\cdot I(A:Y|X)\leq k\cdot\varepsilon\cdot I(X:Y).$
@@
It remains to prove the moreover part of the lemma. We already know that the entropy profile $(X,Y,A,B,C_{0},\ldots,C_{n-1},W_{k})$ is symmetric under swapping $A$ and $B$. To make the property stronger and enforce the symmetry of the tropical tuple, we apply Corollary 4.
3.3 Proof of Corollary E
@@
The argument is based on Lemma D, where we will choose a suitable $k$ and let $W=W_{k}$. We apply Lemma D with the minimal $k$ such that
@@
is at least $\delta\cdot I(X:Y)$. Since the sum of all such gaps does not exceed $I(X:Y)$, the number of steps $k$ is at most $k\leq 1/\delta.$ From Lemma D(c) we get
@@
$I($ $A:X,Y|W_{k})=I(A:X|Y,W_{k})+I(A:Y|X,W_{k})+I(A:X:Y|W_{k})$
$\leq I(A:X|Y)+H(W_{k}|X)+I(A:Y|X)+H(W_{k}|Y)+I(A:X:Y|W_{k})$
$\leq I(A:X|Y)+I(A:Y|X)+I(A:X:Y|W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
$\leq I(A:X:Y|W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
@@ 4 Conditional inequalities with precision $O(\sqrt{\varepsilon})$
First of all, we apply Corollary E and take a $W$ such that
@@
  • $I(A:X,Y|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$.
@@
$\begin{split}I(X:Y|A)&\geq H(W|A)+I(X:Y|A,W)-O(\sqrt{\varepsilon})\cdot I(X:Y)\\ &\geq H(W|A)+I(X:Y|W)-O(\sqrt{\varepsilon})\cdot I(X:Y).\end{split}$
@@
Similarly to the previous argument, we apply Corollary E (switching $A$ and $Y$ in the assumptions and the conclusion of the lemma) and take a $W$ such that
@@
$\begin{split}I(X:Y)&\leq I(X:Y)+I(W:Y|X)\\ &=I(X,W:Y)\\ &=I(W:Y)+I(X:Y|W),\\ &\leq I(W:Y)+I(X:Y|W)+I(A:Y|X,W)\\ &=I(W:Y)+I(X,A:Y|W)\\ &=I(W:Y)+O(\Delta),\end{split}$
@@ 5 Subtler claims with the $O(\sqrt{\varepsilon})$ precision
In this section we combine Lemma D with symmetry considerations, derive a new information inequality for six-tuples of random variables, and prove Theorem F from the introduction.
@@
First of all, we observe that the inequalities in both claims (a) and (b) are symmetric as syntactic expressions with respect to the transposition of $A$ and $B$. Therefore, we can apply Lemma 3 (see also Remark 5) and assume without loss of generality that $A$ and $B$ are symmetric in the distribution $(X,Y,A,B,A^{\prime},B^{\prime})$. Now we are ready to prove the theorem. To address the first claim of the theorem, we apply Corollary E to construct $W$ such that
@@
  • $I(A:X,Y|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$.
Due to Lemma D (the moreover part) we may assume that all entropy quantities for $(X,Y,A,W)$ are the same as the corresponding quantities for $(X,Y,B,W)$. Similarly to (5), using the symmetry we have
@@
On the other hand, for all $(X,Y,A^{\prime},B^{\prime},W)$ we can follow the usual scheme of the proof of the Ingleton’s inequality with $W$ representing the common information of $X$ and $Y$. More formally, we add Shannon–type inequalities
$H(W)$ $\leq H(W|A^{\prime})+H(W|B^{\prime})+I(A^{\prime}:B^{\prime}),$
$H(W|A^{\prime})$ $\leq H(W|X)+H(W|Y)+I(X:Y|A^{\prime}),$
$H(W|B^{\prime})$ $\leq H(W|X)+H(W|Y)+I(X:Y|B^{\prime}),$
@@
$\begin{split}I(X:Y)&\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})\\ &+O(\sqrt{\varepsilon})\cdot I(X:Y)+I(X:Y|W)\end{split}$ (10)
@@
$I(X:Y)\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})+\big(O(\sqrt{\varepsilon})+t\big)\cdot I(X:Y).$
@@ 6 Inequalities with the precision of $O(\varepsilon\log\frac{1}{\varepsilon})$
Now we use Lemma D and symmetry of the distribution to find a sequence of extensions $(X,Y,A,B,W_{i})$, $i=0,1,\dots$ satisfying
@@
$\delta\cdot I(X:Y)\leq I(X:Y|W_{k})\leq\lambda I(X:Y|W_{k-1})\leq\ldots\leq\lambda^{k}\cdot I(X:Y)$
Therefore
@@
We apply inequality (25) from the Appendix, and use the properties of $W_{i}$ from Lemma D to estimate
@@
In addition, we use three other Shannon–type inequalities:
@@
(which is equivalent to the sum of $H(W|X)\geq 0$ and $I(W:X|Y)\geq 0$),
@@
Finally, we use a version of the classical non-Shannon–type inequality from [ZY98] in the form with five variables (proven in [Mak+02]),

### v2

@@ 3.2 Proof of Lemma E
$=I(\tilde{A}_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1})-I(\tilde{A}_{k}:X)$
$=I(\tilde{A}_{k}:X,Y)-I(\tilde{A}_{k}:X)$
/* from the Copy Lemma: $\tilde{A}_{k}$ is independent of
$(A,B,C_{0},\ldots,C_{n-1})$ conditioned on $(X,Y)$ */
@@
$H(W_{k}|X)\leq k\cdot\left(H(A|X)-H(A|X,Y)\right)=k\cdot I(Y:A|X)\leq k\cdot\varepsilon\cdot I(X:Y).$
@@
It remains to prove the moreover part of the lemma. We already know that the entropy profile $(X,Y,A,B,C_{0},\ldots,C_{n-1},W_{k})$ is symmetric under swapping $A$ and $B$. Applying Lemma 6 to the distribution proves the claim.
3.3 Proof of Corollary F
@@
The argument is based on Lemma E, where we will choose a suitable $k$ and let $W=W_{k}$. We apply Lemma E with the minimal $k$ such that
@@
is at least $\delta\cdot I(X:Y)$. Since the sum of the gaps
$\big(I(X:Y)-I(X:Y|W_{1})\big)$ $+\big(I(X:Y|W_{1})-I(X:Y|W_{2})\big)+\ldots$
$\ldots+\big(I(X:Y|W_{k-1})-I(X:Y|W_{k})\big)=I(X:Y)-I(X:Y|W_{k})$
does not exceed $I(X:Y)$, the number of steps $k$ is at most $k\leq 1/\delta.$ From Lemma E(c) we get
@@
$I($ $X,Y:A|W_{k})=I(X:A|Y,W_{k})+I(Y:A|X,W_{k})+I(X:Y:A|W_{k})$
$=I(X:A|Y,W_{k})+I(Y:A|X,W_{k})+I(X:Y:A|W_{k})$
$\leq I(X:A|Y)+H(W_{k}|Y)+I(Y:A|X)+H(W_{k}|X)+I(X:Y:A|W_{k})$
(here we used twice (27) with a suitable substitution of variables)
$\leq I(X:A|Y)+I(Y:A|X)+I(X:Y:A|W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
$\leq I(X:Y:A|W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
@@ 4 Conditional inequalities with precision $O(\sqrt{\varepsilon})$
First of all, we apply Corollary F and take a $W$ such that
@@
  • $I(X,Y:A|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$.
@@
$\begin{split}I(X:Y|A)&\geq H(W|A)+I(X:Y|A,W)-O(\sqrt{\varepsilon})\cdot I(X:Y)\\ &\geq H(W|A)+I(X:Y|W)-I(X:Y:A|W)-O(\sqrt{\varepsilon})\cdot I(X:Y)\\ &\geq H(W|A)+I(X:Y|W)-I(X,Y:A|W)-O(\sqrt{\varepsilon})\cdot I(X:Y)\\ &\geq H(W|A)+I(X:Y|W)-O(\sqrt{\varepsilon})\cdot I(X:Y).\end{split}$
@@
Similarly to the previous argument, we apply Corollary F (switching $A$ and $Y$ in the assumptions and the conclusion of the lemma) and take a $W$ such that
@@
$\begin{split}I(X:Y)&\leq I(X:Y)+I(Y:W|X)\\ &=I(X,W:Y)\\ &=I(Y:W)+I(X:Y|W),\\ &\leq I(Y:W)+I(X:Y|W)+I(Y:A|X,W)\\ &=I(Y:W)+I(X,A:Y|W)\\ &=I(Y:W)+O(\Delta),\end{split}$
@@ 5 Subtler claims with the $O(\sqrt{\varepsilon})$ precision
In this section we combine Lemma E with symmetry considerations, derive a new information inequality for six-tuples of random variables, and prove Theorem D from the introduction.
@@
Remark 15.
The constant hidden in the $O(\cdot)$-notation does not depend on $t$ nor on the distribution of $(X,Y,A,B,C,D)$.
@@
We first show that, without loss of generality, we may assume that the distribution of the tuple $(X,Y,A,B,C,D)$ is symmetric with respect to transposition of $A$ and $B$. We begin with tropicalization: we replace the original tuple by its tropical version $(X,Y,A,B,C,D)^{tr}$, where each variable is replaced by the joint of $n$ independent copies of itself. The assumptions of the theorem still hold for the tropical tuple; see Remark 5. We then apply Lemma 6 and transform this tropical tuple into a symmetric one that still satisfies the assumptions of the theorem. Note that the entropies of the symmetrized tropical tuple coincide with those of the original tuple. Thus, if the theorem holds for the symmetrized tuple, then it also holds for the original one.
Now we are ready to prove the theorem. To address the first claim of the theorem, we apply Corollary F to construct $W$ such that
@@
  • $I(X,Y:A|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$.
Due to Lemma E (the moreover part) we may assume that all entropy quantities for $(X,Y,A,W)$ are the same as the corresponding quantities for $(X,Y,B,W)$. Similarly to (7), using the symmetry we have
@@
On the other hand, for all $(X,Y,C,D,W)$ we can follow the usual scheme of the proof of the Ingleton’s inequality with $W$ representing the common information of $X$ and $Y$. More formally, we add Shannon–type inequalities
$H(W)$ $\leq H(W|C)+H(W|D)+I(C:D),$
$H(W|C)$ $\leq H(W|X)+H(W|Y)+I(X:Y|C),$
$H(W|D)$ $\leq H(W|X)+H(W|Y)+I(X:Y|D),$
@@
$\begin{split}I(X:Y)&\leq I(X:Y|C)+I(X:Y|D)+I(C:D)\\ &+O(\sqrt{\varepsilon})\cdot I(X:Y)+I(X:Y|W)\end{split}$ (12)
@@
$I(X:Y)\leq I(X:Y|C)+I(X:Y|D)+I(C:D)+\big(O(\sqrt{\varepsilon})+t\big)\cdot I(X:Y).$
@@ 6 Inequalities with the precision of $O(\varepsilon\log\frac{1}{\varepsilon})$
The conclusion of the theorem (the Ingleton inequality) is trivial in case $I(X:Y)=0$. So, in what follows we assume without loss of generality that the mutual information $I(X:Y)$ is strictly positive.
@@
Now we use Lemma E and symmetry of the distribution to find a sequence of extensions $(X,Y,A,B,W_{i})$, $i=0,1,\dots$ satisfying
@@
$\delta\cdot I(X:Y)\leq I(X:Y|W_{k})\leq\lambda I(X:Y|W_{k-1})\leq\ldots\leq\lambda^{k}\cdot I(X:Y).$
Since we assume that $I(X:Y)>0$, we may conclude that
@@
We apply inequality (29) from the Appendix, and use the properties of $W_{i}$ from Lemma E to estimate
@@
In addition, we use several other Shannon–type inequalities. First of all, we use
@@
(which is equivalent to the sum of $H(W|X)\geq 0$ and $I(X:W|Y)\geq 0$) and its conditional version
$I(X:Y|Z)\leq H(W|Z)+I(X:Y|W,Z).$ (27)
Besides, we need the inequality
@@
Finally, we use a version of the classical non-Shannon–type inequality from [ZY98] in the form with five variables (proven in [MMRV02, theorem 1 for $n=2$]),
$H(A,B)+2I(X:Y:W)\leq I(X:Y|A)+I(X:Y|B)+H(A)+H(B)+I(XY:W).$
This inequality is equivalent to
$\begin{split}2I(X:Y:W)&\leq I(X:Y|A)+I(X:Y|B)+I(A:B)\\ &\quad+\left(I(X:Y:W)+I(X:W|Y)+I(Y:W|X)\right).\end{split}$
Since $I(X:Y)=I(X:Y:W)+I(X:Y|W)$, it can be rewritten in the form convenient for our application: